### v2

@@ Abstract
Results: We introduce in this paper a novel paradigm (JESTR) for annotation. Unlike prior approaches that explicitly construct molecular fingerprints or spectra, JESTR leverages the insight that molecules and their corresponding spectra are views of the same data and effectively embeds their representations in a joint space. Candidate structures are ranked based on cosine similarity between the embeddings of query spectrum and each candidate. We evaluate JESTR against mol-to-spec and spec-to-FP annotation tools on three datasets. On average, for rank@[1-5], JESTR outperforms other tools by 23.6% - 71.6%. We further demonstrate the strong value of regularization with candidate molecules during training, boosting rank@1 performance by 11.4% and enhancing the model’s ability to discern between target and candidate molecules. Through JESTR, we offer a novel promising avenue towards accurate annotation, therefore unlocking valuable insights into the metabolome.
@@ 1 Introduction
Analyzing biological samples using untargeted metabolomics, where masses of thousands of metabolites within a biological sample are detected, presents unprecedented opportunities to characterize the metabolome. Annotation, the process of assigning chemical structures to metabolomics measurements, however is riddled with uncertainty. Naïvely, one can presume that the measured mass could be used to determine a metabolite’s molecular structure. However, a particular molecular mass can map to possibly thousands of candidate molecular structures that share the same chemical formula (e.g., there are 44,374 known molecular structures associated with $C_{12}H_{18}N_{2}O_{2}$).
@@
Several techniques address the spectra annotation problem. The go-to technique is “spec-to-spec” comparison (Figure 1B) of the measured query spectra against spectra that are cataloged in spectral reference libraries (Kind et al., 2018). However, despite the growth in spectral libraries, e.g., GNPS (Wang et al., 2016), NIST (NIST, 2020), MoNA (MoNA, 2024), annotation rates remain extremely low due to the imitated coverage of spectral libraries in comparison to the space of all potential molecules. In addition, measured spectra vary tremendously under differing instrument settings, e.g., ionization energy, solvent type, and adduct formation (additional functional groups attached or removed from the ionized molecule). A molecule therefore may have many corresponding spectra, which further limits library coverage. A recent search of spectra within 15,327 datasets (Wang et al., 2016) deposited in the MassIVE (Mass Spectrometry Interactive Virtual Environment) database against 586,647 reference spectra cataloged in the GNPS reference library yielded a positive identification rate of 2.3% (Martin et al., 2024) (Figure 1C).
@@
Figure 2: Novelty of the JESTR annotation approach. A. JESTR avoids the explicit generation of spectra, molecules, and fingerprints, and ranks the candidate molecules against the query spectrum based on their joint-space embeddings. B. JESTR learns to place representations of matching molecule-spectrum pairs close in the joint embedding space relative to non-matching pairs. Further, JESTR utilizes additional molecules beyond those in the training set to learn to distinguish target molecules in the training dataset from candidate molecules (those with similar molecular formulas).
We address in this paper the problem of assigning chemical structures from a candidate set to spectral data. The novelty of our approach lies in avoiding the explicit generation of molecular fingerprints and spectra (Figure 2A), and in considering a molecule and its spectra as views of the same object (Figure 2B). This valuable insight allows us to embed molecules and their matching spectra close in the molecule-spectrum joint embedding space. Our approach avoids the need for any kind of reconstruction to intermediate forms such as fingerprints or spectra and therefore removes any reconstruction loss that would invariably creep into the ranking pipeline with any reconstruction based approach. The ranking of candidate structures can be attained by comparing their embeddings against that of the query spectrum and selecting the candidate with the highest cosine similarity. The idea of learning joint embedding spaces from multiple views of the data dates back to the seminal work on Siamese Networks (Chopra et al., 2005). More recently, CLIP (Contrastive Language- Image Pre-training) was trained to create a shared embedding space for both images and text, enabling the model to match relevant images and captions without the need for direct labeling or supervised training on specific dataset (Radford et al., 2021). As we embed molecules and spectra in a joint embedding space, our method is termed Joint Embedding Space Technique for Ranking candidate molecules, JESTR. We use CMC, Contrastive Multiview Coding (Tian et al., 2020), to learn view-invariant information across different views of the data and produce embeddings in a joint embedding space. Recent work in contrastive learning (Chen et al., 2020; Tian et al., 2020; Khosla et al., 2020) has extended the Siamese network approach significantly. Broadly speaking, in all these approaches the key question is how to generate the paired views (either appearing naturally or generated via data augmentation); and how to ensure that paired views end up close together in the joint embedding space.
@@
We conduct experiments and analysis to answer the following research questions. (Q1) Does JESTR’s implicit annotation method (with and without regularization) outperform prior explicit methods (mol-to-spec and spec-to-fp)? (Q2) Is learning the molecule-spectrum joint embedding space effective for distinguishing target molecules among their respective candidate sets? We compare JESTR against state-of-the-art mol-to-spec technique, ESP (Li et al., 2024), and spec-to-FP technique, MIST (Goldman et al., 2023). We conduct the evaluation using three datasets: the NPLIB1 dataset that was previously released with the CANOPUS tool (Dührkop et al., 2021), the well-curated, available-for-purchase NIST2020 dataset, and user-deposited data from MassBank of North America (MoNA) (MoNA, 2024). The contributions of this paper are:
@@
  • Fine-tuning the implicit model via regularization using the candidate sets of the training molecules improves the rank @1 performance in the range of 6.04% to 37.11% when compared to a baseline that does not utilize regularization.
  • Demonstrating that JESTR outperforms ESP and MIST on all ranking metrics and all datasets with the exception of rank@1 for the MoNA dataset. For rank@[1 through 5], JESTR outperforms ESP by 71.6% and MIST by 23.6% across three datasets. These remarkable improvements are achieved even though JESTR does not utilize the additional data in the form of chemical formulae labels for spectra peaks that is currently used by MIST.
@@ 2 Methods
The JESTR model architecture (Figure 3) consists of a molecular encoder and a spectral encoder. They are trained to create embeddings in a molecule-spectrum joint embedding space. To place views of the same object close to each other in the embedding space, we use the CMC contrastive learning loss (Tian et al., 2020). To improve performance, we utilize regularization. At inference, when provided a candidate set for the query spectrum, the cosine similarity is computed between each candidate and the query spectrum. The candidates are then ranked based on their cosine similarities.
@@ 3.1 Datasets
The NPLIB1 dataset was first utilized by the CANOPUS tool to predict compound classes, e.g., benzenoids, phenol ethers, and others, from spectra, thus providing partial annotation on spectra in cases when spec-to-spec comparisons in reference spectral databases yield unsatisfactory matches (Dührkop et al., 2021). This dataset was created by selecting spectra from the NIST2020 (NIST, 2020), GNPS (Wang et al., 2016) and MoNA (MoNA, 2024) databases. The selection ensured a desired distribution of compound classes. This dataset was recently renamed to NPLIB1 (Goldman et al., 2023) to distinguish it from the CANOPUS tool. We utilized the NPLIB1 data as assembled by MIST (Goldman et al., 2023). This dataset comprised 8,030 spectra measured under positive mode (positively charged, with an H adduct, [M+H]+) belonging to 7,131 unique target molecules. We utilize the same data split as proposed by MIST, where the split was structure-disjoint such that a molecule with the same InChiKeys did not appear both in the training and test sets. Therefore, 714 molecules and their 819 spectra were utilized for testing. Two additional datasets were utilized to explore training JESTR on larger datasets. The NIST2020 dataset is well-curated spectral database released by the National Institute of Standards and Technology. NIST2020 comprises a variety of molecules from human, bacteria, environmental, plant, and food samples. A variety of instruments and settings are used to measure spectra for each compound. The measurements are repeated and a consensus spectrum is created for each measurement. The NIST datasets are available under a commercial license, and we had access to the NIST2020 version of this dataset. The MassBank of North America (MoNA) is a collaborative database, with contributions by users. Both experimental and in-silico spectra are accepted. Here, we only retrieved the experimental dataset. Statistics for the three datasets is provided in Table 1. The number of unique molecules is largest in NIST2020, while the number of spectra per molecule is the lowest in NPLIB1. The splits for the NIST2020 and MoNA were created ensuring that no molecules overlapped between the training and test sets.
@@
As typical in prior works (Dührkop et al., 2015; Huber et al., 2021; de Jonge et al., 2023), we select candidates for each target molecule in the training and test data by retrieving molecules from PubChem (Kim et al., 2019), by matching formulae of the target molecules. The average candidate sets for the target molecules range from 1,322 to 2,494 molecules, and for regularization, the average candidate sets for training molecules ranged from a 1,390 to 2,322 (Table 1).
@@
Dataset Spectra Molecules Spectra Molecules Max Cands Avg Cands Spectra Molecules Max Cands Avg Cands
@@
MoNA 35,752 6,767 32,216 6,090 42,542 2,322 677 3,536 32,364 2,494
Table 2: Ranking results for JESTR for the NPLIB1, NIST2020, and MoNA datasets. A. Ranking performance of all three tools B. Ablation study for the impact of removing regularization.
NPLIB1 NIST2020 MoNA
A. Comparison with other tools
Model Rank@1 Rank@5 Rank@20 Rank@1 Rank@5 Rank@20 Rank@1 Rank@5 Rank@20
ESP 23.7 44.7 60.2 20.5 30.1 48.8 19.4 38.4 53.0
MIST 27.9 62.4 82.1 21.0 57.4 78.8 32.3 52.9 75.3
JESTR (w/ regularization) 45.8 81.5 95.8 38.6 59.6 81.0 26.6 64.2 91.2
B. Ablation study
JESTR (w/o regularization) 41.1 81.1 96.1 36.4 56.5 77.9 19.4 52.8 83.9
@@ 3.2 Other annotation tools
To compare with other annotation models, we trained ESP (Li et al., 2024) and MIST (Goldman et al., 2023) on each of the datasets. The ESP model utilizes a GNN-based molecular encoder and an MLP on the molecular fingerprint. ESP is trained to learn a weighting between the molecular and fingerprint representations to predict the spectra. The best ESP performing model on rank@1 was the version that utilized the fingerprint and modeled peak co-dependencies, ESP MLP-PD. MIST first assigns chemical formulas to peaks within each spectrum using SIRIUS (Dührkop et al., 2019), and represents a spectrum as a set of chemical formulas. MIST trains a transformer model to learn peak embeddings and to predicts fingerprint. MIST also featurizes pairwise neutral losses and predicts substructure fragments as an auxiliary task. We ran both MIST and ESP on all three datasets, and confirmed the results with the respective teams.

### v3

@@ Abstract
Results: We introduce in this paper a novel tool (JESTR) for annotation. Unlike prior approaches that explicitly construct molecular fingerprints or spectra, JESTR leverages the insight that molecules and their corresponding spectra are views of the same data and effectively embeds their representations in a joint space. Candidate structures are ranked based on cosine similarity between the embeddings of query spectrum and each candidate. We evaluate JESTR against mol-to-spec, spec-to-FP, and spec-mol matching annotation tools on four datasets. On average, for rank@[1-20], JESTR outperforms other tools by 55.5% - 302.6%. We further demonstrate the strong value of regularization with candidate molecules during training, boosting rank@1 performance by 5.72% across all datasets and enhancing the model’s ability to discern between target and candidate molecules. When comparing JESTR’s performance against that of publicly available pretrained models of SIRIUS and CFM-ID on appropriate subsets of MassSpecGym dataset, JESTR outperforms these tools by 31% and 238%, respectively. Through JESTR, we offer a novel promising avenue towards accurate annotation, therefore unlocking valuable insights into the metabolome.
@@ 1 Introduction
Analyzing biological samples using untargeted metabolomics, where masses of thousands of metabolites within a biological sample are detected, presents unprecedented opportunities to characterize the metabolome. Annotation, the process of assigning chemical structures to metabolomics measurements, however is riddled with uncertainty. Naïvely, one can presume that the measured mass could be used to determine a metabolite’s molecular structure. However, a particular molecular mass can map to possibly thousands of candidate molecular structures that share the same chemical formula. For example, there are 44,374 known molecular structures in PubChem that are associated with chemical formula $C_{12}H_{18}N_{2}O_{2}$). Advanced computational methods promise solutions to address the challenges of annotation uncertainty. IPA (Del Carratore et al., 2019) and ipaPy2 (Del Carratore et al., 2023) enhance confidence in metabolite identification by integrating additional data such as isotope patterns and adduct formation, to provide statistically rigorous probability estimates for each annotation. Mummichog (Li et al., 2013) shifts the focus from individual metabolites to entire metabolic pathways, leveraging peak co-occurrence to infer biological relevance even when exact structures remain unknown. PUMA (Hosseini et al., 2020) further advances annotation by applying probabilistic modeling, predicting pathway activity and assigning chemical identities based on generative inference. Together, these tools improve annotation accuracy, enhance biological interpretation, and enable a more comprehensive understanding of the metabolome. Despite these advances based on analyzing the ionized metabolites, annotation rates remain low.
@@
Several techniques address the spectra annotation problem. The go-to technique is “spec-to-spec” comparison (Figure 1B) of the measured query spectra against spectra that are cataloged in spectral reference libraries (Kind et al., 2018). Searching spectral libraries using learned embeddings, e.g., Spec2Vec (Huber et al., 2021), MSBERT (Zhang et al., 2024), DreaMS (Bushuiev et al., 2024a), can improve search performance. However, despite embedding advances and growth in spectral libraries, e.g., GNPS (Wang et al., 2016), NIST (NIST, 2020), MoNA (MoNA, 2024), annotation rates remain extremely low due to the imitated coverage of spectral libraries in comparison to the space of all potential molecules. In addition, measured spectra vary tremendously under differing instrument settings, e.g., ionization energy, solvent type, and adduct formation (additional functional groups attached or removed from the ionized molecule). A molecule therefore may have many corresponding spectra, which further limits library coverage. A recent search of spectra within 15,327 datasets (Wang et al., 2016) deposited in the MassIVE (Mass Spectrometry Interactive Virtual Environment) database against 586,647 reference spectra cataloged in the GNPS reference library yielded a positive identification rate of 2.3% (Martin et al., 2024) (Figure 1C).
@@
We address in this paper the problem of assigning chemical structures from a candidate set to spectral data. The novelty in our approach lies in avoiding the explicit generation of molecular fingerprints and spectra (Figure 2A), and in considering a molecule and its spectra as views of the same object (Figure 2B). One way of avoiding the explicit construction was recently suggested by ChemEmbed (Faizan-Khan et al., 2025), which trains a spectra encoder to predict the Mol2Vec molecular embeddings (Jaeger et al., 2018), effectively defining the embeddings based on the Mol2Vec embedding space. The molecule-spectra multi-modal view insight allows us to embed molecules and their matching spectra close in the molecule-spectrum joint embedding space. Our approach avoids the need for any kind of reconstruction to intermediate forms such as fingerprints or spectra and therefore removes any reconstruction loss that would invariably creep into the ranking pipeline with any reconstruction based approach. The ranking of candidate structures can be attained by comparing their embeddings against that of the query spectrum and selecting the candidate with the highest cosine similarity. The idea of learning joint embedding spaces from multiple views of the data dates back to the seminal work on Siamese Networks (Chopra et al., 2005). More recently, CLIP (Contrastive Language-Image Pre-training) was trained to create a shared embedding space for both images and text, enabling the model to match relevant images and captions without the need for direct labeling or supervised training on specific dataset (Radford et al., 2021). As we embed molecules and spectra in a joint embedding space, our method is termed Joint Embedding Space Technique for Ranking candidate molecules, JESTR. We use CMC, Contrastive Multiview Coding (Tian et al., 2020), to learn view-invariant information across different views of the data and produce embeddings in a joint embedding space. CMC utilizes Normalized Temperature-scaled Cross-Entropy (InfoNCE) Loss, which incentivizes the model to distinguish true positive pairs from all possible negative pairs in a batch by maximizing the similarity between positive pairs while minimizing similarity with negatives. To assess the closeness of a given pair, CMC relies on a temperature-scaled cosine similarity as a discriminating function. The temperature is a model hyperparameter that controls the sharpness of similarity scores, where lower values sharpens the distinction between positive and negative pairs. In contrast, the recent contrastive-learning based joint-embedding model CMSSP (Chen et al., 2024) applies cross-entropy loss to the dot product of spectral and molecular embeddings after softmax, aiming for high embedding similarity for positives and low embedding similarity for negatives. Broadly speaking, in all contrastive learning approaches, e.g., (Chen et al., 2020; Tian et al., 2020; Khosla et al., 2020), the key question is how to generate the paired views (either appearing naturally or generated via data augmentation); and how to ensure that paired views end up close together in the joint embedding space.
@@
We conduct experiments and analysis to answer the following research questions. (Q1) Does JESTR’s implicit annotation method (with and without regularization) outperform prior explicit methods (mol-to-spec and spec-to-fp)? (Q2) Is learning the molecule-spectrum joint embedding space effective for distinguishing target molecules among their respective candidate sets? We compare JESTR against state-of-the-art mol-to-spec technique, ESP (Li et al., 2024), spec-to-FP technique, MIST (Goldman et al., 2023), and the recent mol-spec matching technique, CMSSP (Chen et al., 2024). We conduct the evaluation using four datasets: the NPLIB1 dataset that was previously released with the CANOPUS tool (Dührkop et al., 2021), the well-curated, available-for-purchase NIST2020 dataset, and user-deposited data from MassBank of North America (MoNA) (MoNA, 2024), and the recently released benchmarking datasets, MassSpecGym dataset (Bushuiev et al., 2024b). Additionally, to ensure a fair evaluation against SIRIUS and CFM-ID, we identify subsets of the MassSpecGym test set that are disjoint from their pretraining data and use these subsets to assess the performance of JESTR relative to these tools.
The contributions of this paper are:
@@
  • Fine-tuning the implicit model via regularization using the candidate sets of the training molecules improves the rank @1 performance in the range of 0.51% to 37.05% when compared to a baseline that does not utilize regularization.
  • Demonstrating that JESTR outperforms ESP and MIST on all ranking metrics and all datasets with the exception of rank@1 for the MoNA dataset. For rank@[1-20], JESTR outperforms ESP by 55.45%, MIST by 56.6%, and CMSSP by 302.56% across four datasets. On carefully selected subsets of the MassSpecGym dataset, JESTR outperforms SIRIUS and CFM-ID on rank@1 by 31% and 238%, respectively. These remarkable improvements are achieved even though JESTR does not utilize the additional data in the form of chemical formulae labels for spectra peaks that are currently used by MIST.
Figure 2: Novelty of the JESTR annotation approach. A. JESTR avoids the explicit generation of spectra, molecules, and fingerprints, and ranks the candidate molecules against the query spectrum based on their joint-space embeddings. B. JESTR learns to place representations of matching molecule-spectrum pairs close in the joint embedding space relative to non-matching pairs. Further, JESTR utilizes additional molecules beyond those in the training set to learn to distinguish target molecules in the training dataset from candidate molecules (those with similar molecular formulas). During ranking, the candidate whose embeddings have the maximum cosine similarity to spectrum embedding is chosen as the highest ranked molecule for that spectrum.
@@ 2 Methods
The JESTR model architecture (Figure 3) consists of a molecular encoder and a spectral encoder. They are trained to create embeddings in a molecule-spectrum joint embedding space. To place views of the same object close to each other in the embedding space, we use the CMC contrastive learning loss (Tian et al., 2020). To improve performance, we utilize regularization. At inference, when provided a candidate set for the query spectrum, the cosine similarity is computed between each candidate and the query spectrum. The candidates are then ranked based on their cosine similarities. Ranking results are reported using rank@k, which is defined as the percentage of target molecules that are ranked at rank $k$ or better.
@@ 3.1 Datasets
Four datasets were used to evaluate JESTR. The NPLIB1 dataset, or the CANOPUS dataset, was first utilized by the CANOPUS tool to predict compound classes, e.g., benzenoids, phenol ethers, and others, from spectra, thus providing partial annotation on spectra in cases when spec-to-spec comparisons in reference spectral databases yield unsatisfactory matches (Dührkop et al., 2021). This dataset was created by selecting spectra from the NIST2020 (NIST, 2020), GNPS (Wang et al., 2016) and MoNA (MoNA, 2024) databases. The selection ensured a desired distribution of compound classes. This dataset was recently renamed to NPLIB1 (Goldman et al., 2023) to distinguish it from the CANOPUS tool. We utilized the NPLIB1 data as assembled by MIST (Goldman et al., 2023). This dataset comprised 8,030 spectra measured under positive mode (positively charged, with an H adduct, [M+H]+) belonging to 7,131 unique target molecules. We utilize the same data split as proposed by MIST, where the split was structure-disjoint such that a molecule with the same InChiKeys did not appear both in the training and test sets. Therefore, 714 molecules and their 819 spectra were utilized for testing. Two additional datasets were utilized to explore training JESTR on larger datasets. The NIST2020 dataset is well-curated spectral database released by the National Institute of Standards and Technology. NIST2020 comprises a variety of molecules from human, bacteria, environmental, plant, and food samples. A variety of instruments and settings are used to measure spectra for each compound. The measurements are repeated and a consensus spectrum is created for each measurement. The NIST datasets are available under a commercial license, and we had access to the NIST2020 version of this dataset. The MassBank of North America (MoNA) is a collaborative database, with contributions by users. Both experimental and in-silico spectra are accepted. Here, we only retrieved the experimental dataset. Statistics for the three datasets is provided in Table 1. The number of unique molecules is largest in NIST2020, while the number of spectra per molecule is the lowest in NPLIB1. The splits for the NIST2020 and MoNA were created ensuring that no molecules overlapped between the training and test sets.
The fourth dataset is the recent MassSpecGym (Bushuiev et al., 2024b) dataset, which was developed as a benchmark set for candidate ranking, spectra simulation and molecular generation tasks. The spectra consisted of a mix of H and Na adducts, with a total of 231,104 spectra for 32,010 molecules. of high-quality labeled MS/MS spectra collected from public repositories such as MoNA, MassBank, and GNPS, as well as newly measured in-house data. The dataset includes 231,000 spectra corresponding to 29,000 unique molecular structures, making it the largest publicly available MS/MS benchmark. To ensure rigorous evaluation and prevent data leakage, the dataset is split using Maximum Common Edge Subgraph (MCES) distance, which clusters structurally similar molecules together in the same fold. Unlike InChIKey-based splits, this approach ensures that molecules in the test set have a significant structural difference from those in the training set. For the test split, an MCES distance threshold of > 10 was used. Hence, the MassSpecGym test set is more challenging than all other test splits.
@@
For all datasets except MassSpecGym, we select candidates for each target molecule in the training and test data by retrieving molecules from PubChem (Kim et al., 2019), by matching formulae of the target molecules. The average candidate sets for the target molecules range from 1,322 to 2,494 molecules, and for regularization, the average candidate sets for training molecules ranged from a 1,390 to 2,322 (Table 1). For the MassSpecGym dataset, we utilize the candidate sets provided with the released benchmark. Two types of candidates are provided: either based on mass, or on formula of the precursor ion. The maximum number of candidates for each test spectrum is 256.
@@
Maximum # Average # Maximum # Average #
Dataset Spectra Molecules Spectra Molecules of candidates of candidates Spectra Molecules of candidates of candidates
@@
MoNA 35,752 6,767 32,216 6,090 42,542 2,322 3,536 677 32,364 2,494
MassSpecGym 231,104 32,010 194,119 28,840 48,184 2,720 17,556 3,170 256 185
[…]
Explicit-construction approaches Implicit/Matching approaches
[…]
NPLIB1 @1 23.69 27.96 13.92 54.09 45.76 41.06
@5 44.69 62.39 32.97 67.40 81.53 81.12
@20 60.20 82.05 49.08 74.48 95.77 96.13
NIST2020 @1 20.47 20.95 1.90 2.45 38.62 36.38
@5 30.12 57.36 3.95 4.01 59.64 56.45
@20 48.76 78.83 5.34 4.86 81.03 77.94
MoNA @1 19.37 32.30 3.44 7.64 26.56 19.38
@5 38.37 54.93 7.13 11.49 64.21 52.82
@20 53.04 75.28 10.03 12.82 91.16 83.88
MassSpecGym Candidates by Mass @1 10.71 14.64 2.60 3.85 15.13 15.62
@5 24.84 34.87 5.50 5.44 36.75 37.47
@20 42.66 59.15 8.70 7.71 60.32 60.55
MassSpecGym Candidates by formula @1 11.05 9.57 2.30 3.61 11.85 11.82
@5 27.42 22.11 5.10 5.66 32.95 33.48
@20 52.20 41.10 8.70 8.53 61.46 61.46
@@ 3.2 Other annotation tools
We selected three recent annotation tools to provide a comparative evaluation for JESTR: ESP (Li et al., 2024), MIST (Goldman et al., 2023) and CMSSP (Chen et al., 2024). ESP and MIST are recent representative methods in the mol-to-spec and spec-to-mol categories, respectively, following the approach of explicit generation of intermediate forms, while CMSSP is the only model other than JESTR that matches spectrum and candidates directly in the embedding space. The ESP model utilizes a GNN-based molecular encoder and an MLP on the molecular fingerprint. ESP is trained to learn a weighting between the molecular and fingerprint representations to predict the spectra. The best ESP performing model on rank@1 was the version that utilized the fingerprint and modeled peak co-dependencies, ESP MLP-PD. MIST first assigns chemical formulas to peaks within each spectrum using SIRIUS (Dührkop et al., 2019), and represents a spectrum as a set of chemical formulas. MIST trains a transformer model to learn peak embeddings and to predicts fingerprint. MIST also featurizes pairwise neutral losses and predicts substructure fragments as an auxiliary task. CMSSP is similar to JESTR in utilizing contrastive learning in a joint embedding space, but utilizes a discriminator function based on the dot product.
To streamline the comparison using the same data splits for all datasets, we trained MIST and ESP on all datasets, and confirmed the results with the respective teams. We evaluated CMSSP on all test datasets using CMSSP’s pretrained released weights. The released CMSSP was previously trained on MS/MS spectra from two public databases, GNPS and MassBank, and was supplemented by 1,906 spectra independently acquired in-house. To ensure that there is no data leakage between the training set and test set molecules, we also publish results by retraining CMSSP on the same training-test split as was used when reporting performance for the other models. ESP was trained for 100 epochs with a learning rate of 0.001 and batch size of 32. Adam was used as the optimizer with L2 norm of $10^{-6}$. MIST was trained for 500 epochs with a learning rate of 0.00057 and batch size of 32. For CMSSP, we evaluate the model using the released pretrained weights and also using retrained models on the specific datasets.